OpenStudy (anonymous):

What is the equation of the line passing through (–3, 6) and parallel to the line y = 4x – 2 in standard form? x – 4y = –18 x + 4y = 18 4x – y = –18 4x + y = 18

OpenStudy (anonymous):

plug for y1 and x1 y - y1 = m (x - x1) m=4

OpenStudy (anonymous):

y-6=4(x+3) y-6=4x+12 +6 +6 y=4x+18 -4x -4x -4x+y=18

OpenStudy (anonymous):

so would that be 4x-y=-18

OpenStudy (anonymous):

yes if you multy. by -1

OpenStudy (anonymous):

thank youuu

OpenStudy (anonymous):

YW !